# Bid Farewell to Seesaw: Towards Accurate Long-tail Session-based Recommendation via Dual Constraints of Hybrid Intents

### v1

@@ D.2   Computation Resources
  • STAMP (liu18Stamp) explores the capability of attention layers on session-based recommendation instead of RNNs. It optimizes the attention mechanism of previous work by emphasizing the user’s short-term memory.
  • GRU4Rec (Hidasi16gru4rec) is An RNN based deep learning model for session based recommendation, which utilizes session-parallel mini-batch training process and also employs ranking-based loss functions during the training.
  • SR-GNN (Wu19Srgnn) employs GNNs to learn item embeddings and fuse the item-level information to get the session representation by leveraging the soft-attention mechanism.
  • GCE-GNN (Wang20Gcegnn) constructs two types of graphs to capture the global and local information from input sessions and combine them to enhance the feature presentations of items.
@@
  • TailNet (Liu2020TailNet) is the first classical work in SBR to consider recommendation diversity through the preference mechanism to adjust the importance of tail and head items.
  • CSBR (Chen2023CSBR) addresses the long-tail issue of recommendations with two additional training objectives including the distribution prediction and distribution alignment.
  • LOAM (Yang2023LOAM) address the long-tail issue of recommendation results through the niche-walk augmentation and the tail session mixup.
  • LAP-SR (Peng2024LAPSR) is a post-processing approach that aims to alleviate the long-tail impact in session-based recommender systems by using personalized diversity.
Since HID focuses on addressing the long-tail issue, existing intent-based SBR models (Chen22Intent; Choi2024Intent; Zhang23Efficiently; Wang2024HearInt) which only concentrate on the accuracy of recommendations does not serve as the competitors. Besides, Since MELT (Kim2023MELT) , GUME (Lin2024GUME), GALORE (Luo2023Improving), and LLM-ESR (Liu24LLMESR) utilize collaborative signals from users, which are not available in session-based recommendation due to the anonymity, we have not included them in the competitors either.
@@
To evaluate the recommendation accuracy and long-tail performance, we employ three widely used accuracy metrics, which are HR@K, and MRR@K. Following previous works on long-tail issue (Himan19Managing; Liu2020TailNet; Yang2023LOAM), we introduce some long-tail metrics which are tHR@K, tMRR@K, tCov@K, and Tail@K. tNDCG@K, HRt@K, and MRRt@K calculate the Normalized Discounted Cumulative Gain, Hit Ratio, and Mean Reciprocal Rank of sessions whose next-item (i.e, ground turth item) belongs to the tail items. For the other two long-tail metrics, we give clear definitions as follows:
tCov@K (Tail Coverage) (Liu2020TailNet; Yang2023LOAM) measures how many different tail items ever appear in the top-K recommendations, which can be formulated as:  $tCov@K=\frac{|\cup_{u\in U}L^{T}_{K}(u)|}{|V|}$, where $L^{T}_{K}(u)$ is the set of long tail items within the top-K recommendations of session $u$.
Tail@K (Liu2020TailNet; Yang2023LOAM) measures how many long-tail items in the top-K for each recommendation list. This metric can be formulated as: $Tail@K=\frac{1}{|U|}\sum_{u\in U}\frac{|L_{K}^{T}(u)|}{K}$, where $U$ is the set of sessions.

### v2

@@ D.2   Computation Resources
  • STAMP (Liu et al. 2018) explores the capability of attention layers on session-based recommendation instead of RNNs. It optimizes the attention mechanism of previous work by emphasizing the user’s short-term memory.
  • GRU4Rec (Hidasi et al. 2016) is An RNN based deep learning model for session based recommendation, which utilizes session-parallel mini-batch training process and also employs ranking-based loss functions during the training.
  • SR-GNN (Wu et al. 2019) employs GNNs to learn item embeddings and fuse the item-level information to get the session representation by leveraging the soft-attention mechanism.
  • GCE-GNN (Wang et al. 2020) constructs two types of graphs to capture the global and local information from input sessions and combine them to enhance the feature presentations of items.
@@
  • TailNet (Liu and Zheng 2020a) is the first classical work in SBR to consider recommendation diversity through the preference mechanism to adjust the importance of tail and head items.
  • CSBR (Chen et al. 2023) addresses the long-tail issue of recommendations with two additional training objectives including the distribution prediction and distribution alignment.
  • LOAM (Yang et al. 2023) address the long-tail issue of recommendation results through the niche-walk augmentation and the tail session mixup.
  • LAP-SR (Peng and Zhou 2024) is a post-processing approach that aims to alleviate the long-tail impact in session-based recommender systems by using personalized diversity.
Since HID focuses on addressing the long-tail issue, existing intent-based SBR models (Chen et al. 2022; Choi et al. 2024a; Zhang et al. 2023; Wang et al. 2024) which only concentrate on the accuracy of recommendations does not serve as the competitors. Besides, Since MELT (Kim et al. 2023) , GUME (Lin et al. 2024), GALORE (Luo et al. 2023), and LLM-ESR (Liu et al. 2024) utilize collaborative signals from users, which are not available in session-based recommendation due to the anonymity, we have not included them in the competitors either.
@@
To evaluate the recommendation accuracy and long-tail performance, we employ three widely used accuracy metrics, which are HR@K, and MRR@K. Following previous works on long-tail issue (Abdollahpouri et al. 2019; Liu and Zheng 2020a; Yang et al. 2023), we introduce some long-tail metrics which are tHR@K, tMRR@K, tCov@K, and Tail@K. tNDCG@K, HRt@K, and MRRt@K calculate the Normalized Discounted Cumulative Gain, Hit Ratio, and Mean Reciprocal Rank of sessions whose next-item (i.e, ground turth item) belongs to the tail items. For the other two long-tail metrics, we give clear definitions as follows:
tCov@K (Tail Coverage) (Liu and Zheng 2020a; Yang et al. 2023) measures how many different tail items ever appear in the top-K recommendations, which can be formulated as:  $tCov@K=\frac{|\cup_{u\in U}L^{T}_{K}(u)|}{|V|}$, where $L^{T}_{K}(u)$ is the set of long tail items within the top-K recommendations of session $u$.
Tail@K (Liu and Zheng 2020a; Yang et al. 2023) measures how many long-tail items in the top-K for each recommendation list. This metric can be formulated as: $Tail@K=\frac{1}{|U|}\sum_{u\in U}\frac{|L_{K}^{T}(u)|}{K}$, where $U$ is the set of sessions.